solve this equation: x/x-5= -2x/x

please and thank you!

Answer by mangopeeler07(462)

x/x-5= -2x/x

Cross multiply
-2x(x-5)=x^2

Distribute
-2x^2+10x=x^2

Divide both sides by x

-2x+10=x

Add 2x to both sides
10=3x

Divide both sides by 3
10/3=x